solve for x and y
(5+2i)+(x+yi)=4-3i

Answer by MissionPossible(18)

(5+2i)+(x+yi)=4-3i
5+x = 4
then x = 4-5 = -1
2i + yi = -3i
then 2 + y = -3
y = -3-2 = -5